angelcook2007:

The function f(x) = −x2 + 28x − 192 models the hourly profit, in dollars, a shop makes for selling sodas, where x is the number of sodas sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? Part B: Determine the x-intercepts. What do these values mean in the context of the problem?

jhonyy9:

A. do you know the way how you find the vertex of a quadratic ?

jhonyy9:

vertex (h,k) \[h = - \frac{ b }{ 2a }\] k = f(h)

angelcook2007:

what does that mean?

Extrinix:

\(h=-\dfrac{\color{lime}{b}}{2\color{cyan}{a}}\) is the equation to find the `x-coordinate` of a quadratic equation, a quadratic equation looks like this: \(\color{cyan}{a}\color{tomato}{x}^2+\color{lime}{b}\color{tomato}{x}+\color{yellow}{c}=0\) From your equation find \(a~b\) and \(c\) \(\color{cyan}{-1}\color{tomato}{x}^2+\color{lime}{28}\color{tomato}{x}-\color{yellow}{192}=f(x)\) \(\color{cyan}{a}=-1\) \(\color{lime}{b}=28\) \(\color{yellow}{c}=-192\) Plug it into or equation \(h=-\dfrac{\color{lime}{28}}{2\color{cyan}{-1}}\) Now just solve that and you'll have your answer.

angelcook2007:

wait, what happened to 192

angelcook2007:

-192 i meant